OpenStudy (anonymous):

How do I solve for r a little confused i got here just need help finishing 1.00062441492=1+r/12

OpenStudy (anonymous):

1.00062441492=1+r/12 First you will multiply 12 by 12 to cancel it out. You will also have to multiply the other side by 12 also. 12*1.00062441492=1+r/12*12 Then once you do that you get: 12.01=1+r Then you will subtract 1 from both sides. 12.01-1=1-1+r Then you will get your final answer of: r=11.01

OpenStudy (anonymous):

Thank you so much i appreciate this its the last part of a problem from a long assignment

OpenStudy (anonymous):

You're welcome, anytime.